OpenStudy (anonymous):

Differentiate a func with wrt another ? f(x) = arcsin((2x)/(1+x^2)) and g(x) = arccos((1-x^2)/(1+x^2)) , then df(x)/dg(x) = ? How do I do this ? df(x)/dx * dx/dg(x) ?

OpenStudy (anonymous):

\[f(x)=\sin ^{-1} ( (2x) / (1+x^2) ) \]

OpenStudy (anonymous):

\[g(x)=\cos ^{-1} ( (1-x^2) / (1+x^2) ) \]

OpenStudy (anonymous):

but seems to lengthy to do. :( ?

OpenStudy (anonymous):

df(x)/dx=2/(1+x^2) and dg(x)/dx=2/(1+x^2) dx/dg(x)=(1+x^2)/2 so df(x)/dg(x)=1

OpenStudy (anonymous):

akanksha: what you mentioned before is right, but (df(x)/dx)/ (dg(x)/dx) would be a better choice!

OpenStudy (anonymous):

Ok right, you did the same thing. There is no other way to do it! :)

OpenStudy (anonymous):

@Abdul , how did y ou do that ??

OpenStudy (anonymous):

OK, you have another way to do it. \[x = \tan \theta\] Thus \[f(x) = \sin^{-1}\sin(2 \theta)\]\[= 2 \theta\] \[g(x) = \cos^{-1} \cos(2 \theta)\]\[= 2 \theta\] so the answer is 1.

OpenStudy (anonymous):

Cool ! Thanks.